how do i solve by substitution?
2x+3y=6
y=2x+2

Answer by jim_thompson5910(35256)

Hint:

2x+3y=6

2x+3(2x+2)=6

2x+6x+6=6

Solve for x. Then use this to find y (use y = 2x+2)